OpenStudy (calculusxy):

Solve by using elimination. -6 = -6y - 12x -8y = 10x - 24 I got (-3.5,8)

OpenStudy (mathmale):

Hello. We worked together about an hour ago. But then it seemed that you just disappeared without a 'thank you' or 'good bye.' What happened? Which method did you use to solve this problem?

OpenStudy (calculusxy):

I am sorry. My internet kept on crashing on me, so I just left it. But thank you so much for your help. I used elimination as the question asked for me to do.

OpenStudy (blackstreet23):

6y = -12x + 6 8y = 10x - 24 8(6y = -12x + 6) -6(8y = 10x - 24) 48y = -96x + 48 -48y = -60x + 144 ?

OpenStudy (mathmale):

blackstreet's approach is right on target, so I will leave you two to work together.

OpenStudy (calculusxy):

my equations were -6 + 12x = -6y -24 + 10x = -8y

OpenStudy (blackstreet23):

yeah i just moved them around

OpenStudy (calculusxy):

(-6 + 12x = -6y)-5 (-24 + 10x = -8y)6 30 - 60x = 30y -144 + 60x = -48y -114 = -18y 8 = y

OpenStudy (blackstreet23):

ohh alright seems you know how to do it lol

OpenStudy (calculusxy):

did you get x= -3.5?

OpenStudy (blackstreet23):

48y = -96x + 48 -48y = -60x + 144 (96-60) x + 144 + 48 = 0 36x = -192 x = 5.333

OpenStudy (blackstreet23):

x = -5.3333333

OpenStudy (calculusxy):

ok i will check out what my mistake was. thank you :)

OpenStudy (blackstreet23):

No problem :)